Two charges lie in a line along the x axis. Charge 1 is q1 0.85 C harge 2 is q2 = 2.05 C. They are each a distance of d 0.095 m from the origin 0,0 What is the distance on the x-axis from the origin at which the electric field will be zero. Give your answer in meters. cosO tan() π| sin cotan0asin0 acos0 atan)acotan( cosh) tanh cotanh) END Degrees O Radians BACKSPACE | DEL| CLEAR Submit Hint Feedback I give up!
Answer #1

Suppose at distance d from charge q1 net electric field is zero, then

Electric field is given by:

E = kQ/R^2

Since we know that direction of electric field due to +ve charge will be away from charge and due to -ve charge will be towards the charge. In given case since both charge is positive, So

due to charge q2 electric field will be toward +ve x-axis and due to charge q1 electric field will be toward -ve x-axis. So

Enet = E2 - E1 = 0

E1 = E2

kq1/r1^2 = kq2/r2^2

q1 = 0.85 C

q2 = 2.05 C

r1 = d = distance from charge q1

r2 = distance from charge q2 = 2*0.095 - d = 0.19 - d

(We used 2*0.095 because total distance between both charges is twice of 0.095 m)

Using above values:

0.85/d^2 = 2.05/(0.19 - d)^2

(0.19 - d)/d = sqrt (2.05/0.85) = 1.5530 m

0.19 - d = 1.5530*d

d = 0.19/(1 + 1.5530)

d = 0.074 m = distance from charge q1

distance from origin will be = 0.095 - 0.074 = 0.021 m

See that we need final distance from origin.

Please Upvote.

Comment below if you have any query.